How much work must be done in blowing a spherical soap bubble 20 cm in diameter? The area of a sphere is 4πr² and the surface tension of the soap solution is 0.032 N/m.
1
Expert's answer
2020-11-30T14:55:18-0500

The work is


"W=A\sigma=4\pi r^2\sigma=d^2\pi\sigma=4.02\text{ mJ}."